1.Ten years ago Diana Torres wrote what has become the leading Tort textbook. She has been receiving royalties based on revenues reported by the publisher. These revenues started at £1,700 in the first year, and grew steadily by 4.9% per year. Her royalty rate is 18% of revenue. Recently, she hired an auditor who discovered that the publisher had been underreporting revenues. The book had actually earned 10% more in revenues than had been reported on her royalty statements. Assuming the publisher pays an interest rate of 4.7% on missed payments, how much money does the publisher owe Diana?



1
Expert's answer
2022-03-13T18:55:33-0400

In the first year, the publisher paid:

"1700\times0.18=306"

taking into account the increase in income:

"((1700\times0.1)+1700)\times0.18=336.6"

in the second year, the publisher had to pay:

"((1700\times0.1)+1700)\times0.18=336.6"

For two years:

336.6+336.6=673.2

in the second year, the publisher paid:

"((1700\times0.049)+1700)\times0.18=320.99"

For two years paid:

306+320.99=626.99

Difference in:

673.2-626.99=57.21